Test the hypothesis using the​ P-value approach. Be sure to verify the requirements of the test.

Upper H 0H0​:

pequals=0.40.4

versus

Upper H 1H1​:

pgreater than>0.40.4

nequals=125125​;

xequals=5555​,

alphaαequals=0.050.05

Is

np 0 left parenthesis 1 minus p 0 right parenthesisnp01−p0greater than or equals≥​10?

No

Yes

Use technology to find the​ P-value.

​P-valueequals=nothing

​(Round to three decimal places as​ needed.)

Do not reject

Reject

the null​ hypothesis, because the​ P-value is

less

greater

than

alphaα.
Answer #1

We have n = 125, x = 55,

The hypothesis are:

H0: P = 0.4 v/s H1: P > 0.4

Here nP = 125*0.4 = 50 > 10, n(1-P) = 125 * 0.6 = 75 > 10

The test statistic is,

= 0.91

p value = p ( z > 0.91 )

= 1 - p ( z 0.91 )

= 1 - 0.8186 -----------( using excel formula "=norm.s.dist(0.91,1)" )

= 0.1814

Here p value >

Hence we do not reject null hypothesis.